# On the Trade-off between Flatness and Optimization in Distributed Learning

### v1

@@ III-A Modeling conditions
We introduce the following commonly-used assumptions. First, since our purpose is to understand the behavior of the algorithm around the same local minima $w^{\star}$, we assume all agents start from points near $w^{\star}$.
@@
(Starting points of all agents). All models start their updates from initial points that are sufficiently close to $w^{\star}$, namely,
@@
By using Jensen’s inequality on (12), we have
@@
Since we will establish that mini-batch gradient descent algorithms approach an $O(\frac{\mu}{B})$ neighborhood of $w^{\star}$, it is reasonable to assume all agents start from the $o(\frac{\mu}{B})$ neighborhood, which is dominated by $O(\frac{\mu}{B})$. This assumption is weaker than the one used in [34] where models should exactly start from $w^{\star}$.
@@
$\bar{H}\overset{\Delta}{=}\frac{1}{K}\sum_{k=1}^{K}\nabla^{2}J_{k}(w^{\star})=\frac{1}{K}\sum_{k=1}^{K}H_{k}(w^{\star})$ (16)
@@ III-B Properties of gradient noise
where the term stochastic refers to the case $B=1$ (i.e., batch size equal to 1). In the mini-batch case, the gradient noise is instead given by:
@@
We can further relate the noise covariance matrix at $w^{\star}$ and the Hessian matrix $\bar{H}$. In the single-agent case, when $B=1$ and negative log-likelihood losses are used, it is well-known that there is an exact equivalence between the Hessian and the gradient covariance matrices at local minima in stochastic risk minimization [34], and this relationship can also be approximately used in the context of empirical risk minimization [34, 53, 39]. By log-likelihood losses we mean choosing $Q_{k}(w;\boldsymbol{x}_{k})$ to be of the form
$Q_{k}(w;\boldsymbol{x}_{k})=-\log p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k};w)$ (32)
where $\boldsymbol{h}_{k}$ and $\boldsymbol{\gamma}_{k}$ represent the feature vector and label associated with $\boldsymbol{x}_{k}$. In the multi-agent case, we define the gradient covariance matrix of the global risk function at $w^{\star}$ for the case $B=1$ as:
@@
Lemma III.5.
(Relation to Hessian matrix). Under Assumption II.1, if the negative log likelihood losses are used by all agents, it holds that
$\bar{R}_{s}\approx\frac{1}{K}\bar{H}$ (34)
Proof.
See Appendix B. ∎
Lemma III.5 establishes the approximate equivalence between $\bar{R}_{s}$ and $\frac{1}{K}\bar{H}$ in the context of empirical risk minimization. Since negative log-likelihood losses, e.g., mean-square loss and cross-entropy loss, are popular in the machine learning community, we mainly focus on this category in this paper, but our framework can also be applied to other losses.
@@ III-C Network performance analysis
This approximation naturally raises the following question: How accurate can the short-term model in (48) approximate the true recursion in (42)? This question can be answered by the following two lemmas, where we separately show the results for the decentralized and centralized methods over a finite time horizon.
Lemma III.6.
@@
See Appendices C, D and E. ∎
Lemma III.7.
(Deviation bounds of the centralized method). Under the same conditions of Lemma III.6, the second-order and fourth-order moments of ${\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}$, the second-order moment of ${\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}$, and the approximation error of the short term model, related to the centralized method are guaranteed to be upper bounded in a finite number of iterations $n\leq O(\frac{1}{\mu})$. Basically, it holds that
@@
$\Big{|}\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}\Big{|}\leq O(\mu^{1.5(1+\eta)})$ (57)
@@
See Appendices C, D and E. ∎
Comparing the bounds in Lemmas III.6 and III.7 , we find that the upper bounds on the second-order moments associated with the decentralized methods incorporate extra $O(\mu^{2})$ terms, which are generated by the network heterogeneity and graph structure. We will discuss the effects of the $O(\mu^{2})$ terms on the escaping efficiency later. Moreover, Lemmas III.6 and III.7 demonstrate that $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}$ and $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}\|^{2}$ dominate $|\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}|$ for all methods when $\mu$ is sufficiently small. In other words, the approximation error between the mean square deviation of the short-term and true models can be omitted compared with the size of the true models. Furthermore, we can manipulate the bounds in Lemmas III.6 and III.7 to obtain expression for $\mathrm{ER}_{n}$ performance. Using (7) and (10), we can approximate the $\mathrm{ER}_{n}$ as follows:
@@
which means that $\mathrm{ER}_{n}$ can be evaluated by means of the short-term recursion. Thus, the original recursion in (42) can be replaced by (48) in the small step-size regime. More details about the equality (III-C) can be found in Appendix F.
Building upon the results of Lemmas III.6 and III.7, we next analyze the escaping efficiency of the three algorithms by using (48), and establish the following theorem.
Theorem III.8.
@@
$e(n)=\frac{1}{4K}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{H}\right)$ (63)
@@
See Appendix F. ∎
Theorem III.8 shows the escaping efficiency of the three algorithms around a local minimum over a finite time horizon. Basically, for all three methods, larger $\mu$ or smaller batch size $B$ (i.e., larger gradient noise) enable higher escaping efficiency. Also, as mentioned in Appendix F, for the algorithms to successfully exit the local basin they would require $O(\frac{1}{\mu})$ iterations. Thus, if the algorithms cannot leave the local minimum in $O(\frac{1}{\mu})$ iterations, then we say they cannot effectively escape the local basin and are therefore trapped in it. Furthermore, we observe that the network heterogeneity and graph structure implicitly introduce additional $O(\mu^{2})$ terms into decentralized methods. That is, for homogeneous networks where all agents share the same local minimum $w^{\star}$, we have
@@
However, the difference between the centralized and decentralized methods is only significant in the large-batch regime. Specifically, if $B$ is not sufficiently large such that $\eta<1$, then we have $21+\eta<$ 2 under which the additional $O(\mu^{2})$ terms in (61) and (62) will be dominated by the $o(\mu^{1+\eta})$ terms associated with the approximated error due to (48) and the initialization condition in Assumption III.1. However, as the batch size $B$ increases, the effect of the gradient noise will progressively decrease and the influence of the $O(\mu^{2})$ terms will become more prominent. In the extreme scenario when the full-batch gradient descent is applied, the centralized method becomes completely noise-free, which means that there is no noise to help the algorithm escape from local minima. Nevertheless, the noisy terms related to network heterogeneity and graph structure continue to exist in decentralized approaches.
When comparing the escaping efficiency of two methods from a local minimum, such as the centralized and diffusion methods, there are three possible cases on average: (1) Both $\mathrm{ER}_{n,cen}$ and $\mathrm{ER}_{n,dif}$ exceed the risk barrier $h$, then the algorithms will leave the basin around the local minimum; (2) $\mathrm{ER}_{n,dif}$ is lager than $h$, while $\mathrm{ER}_{n,cen}$ is smaller than $h$. In this case, the network driven by diffusion escapes from the local minimum, while the centralized algorithm remains stuck in the current basin; (3) Both $\mathrm{ER}_{n,cen}$ and $\mathrm{ER}_{n,dif}$ are smaller than $h$, then the two methods remain trapped in the current basin. These 3 situations illustrate how higher escaping efficiency corresponds to a higher likelihood of escape from a local minimum. It is also clear that a decentralized method is more effective at escaping from a local minimum than the centralized method in the large-batch regime. However, larger values of $\mathrm{ER}_{n}$, while good for escaping efficiency, they nevertheless worsen the optimization performance. We will discuss this point later in the steady state when $n\to\infty$.
@@
Corollary III.9.
(Smaller $\mathrm{ER}_{n}$ for diffusion). Under the same conditions of Theorem III.8, it holds that
@@
See Appendix G. ∎
Corollary III.9 implies that consensus runs farther away from the local minimum than the diffusion strategy for the same number of iterations due to its worse ER performance. This translates into faster escape for consensus compared to diffusion.
@@ IV Trade-off between flatness and optimization
Suppose we run a decentralized or centralized algorithm to solve a possibly nonconvex optimization problem. One useful question is to investigate where the algorithm would prefer to go if it escapes from the current basin. In other words, assume the algorithm escapes from the basin around some local minimum $w^{\star}$ and starts evolving until it settles down in the basin of another local minimum, we would like to examine what preferential properties does this second minimum have relative to the earlier one. To answer this question, we will relate the escaping efficiency of algorithms measured by $\mathrm{ER}_{n}$ and the flatness of local minima measured by $\mathrm{Tr}(\bar{H})$.
@@
$e(n)=$ $\frac{1}{4K}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{H}\right)$
$=$ $\frac{1}{4K}\sum\limits_{i=1}^{M}\lambda_{i}(\bar{H})\left(1-(1-\mu\lambda_{i}(\bar{H}))^{2(n+1)}\right)$
$\leq$ $\frac{1}{4K}\sum\limits_{i=1}^{M}\lambda_{i}(\bar{H})=\frac{1}{4K}\mathrm{Tr}(\bar{H})=\mathcal{U}\big{(}e(n)\big{)}$ (69)
where $\lambda_{i}(\bar{H})$ represents the $i_{th}$ eigenvalue of $\bar{H}$. Also,
@@
Observe that these upper bounds on the escaping efficiency are positively correlated with the sharpness of the local minimum through $\mathrm{Tr}(\bar{H})$. Substituting (72)–(74) into (68) separately, we find that flat minima where $\mathrm{Tr}(\bar{H})$ is small provide high probability guarantees that the algorithms stay around the corresponding local basin.
Combining the discussion of both one and higher-dimensional cases, we conclude that the three algorithms prefer to stay around flat minima, which means that they favor flatter basin if they successfully escape from a current minimum. Moreover, as discussed before, higher escaping efficiency makes decentralized algorithms more likely to escape from a local minimum than the centralized method. This fact further indicates that decentralized methods favor flatter minima than their centralized counterpart. However, as already noted before, higher escaping efficiency may deteriorate the optimization performance. This motivates us to analyze the excess-risk performance in the long run when $n\to\infty$, which corresponds to the optimization performance in the steady state.
@@
(Steady-state excess risks). Consider a network of agents running distributed algorithms covered by (48). After sufficient iterations such that $n\to\infty$, under assumptions II.1, III.1, III.2 and III.3, and assuming the algorithms are already trapped in the basin of a local minimum $w^{\star}$ and $J(w)$ is locally strongly-convex around $w^{\star}$, it holds that
$\mathrm{ER}_{\infty,cen}=$ $\frac{\mu}{4BK}\mathrm{Tr}\left(\bar{H}\right)\pm O\left(\mu^{1.5(1+\eta)}\right)$ (75)
$\mathrm{ER}_{\infty,con}=$ $\frac{\mu}{4BK}\mathrm{Tr}\left(\bar{H}\right)+\frac{\mu^{2}}{2K}\|d^{\sf T}\mathcal{V}_{\alpha}(I-\mathcal{P}_{\alpha})^{-1}\|^{2}_{I\otimes\bar{H}}$
@@
$\mathrm{ER}_{\infty,dif}=$ $\frac{\mu}{4BK}\mathrm{Tr}\left(\bar{H}\right)+\frac{\mu^{2}}{2K}\|d^{\sf T}\mathcal{V}_{\alpha}P_{\alpha}(I-\mathcal{P}_{\alpha})^{-1}\|^{2}_{I\otimes\bar{H}}$
@@
See Appendix H. ∎
In Corollary IV.1, the factors that help algorithms escape local minima are now seen to adversely affect the optimization performance. Again, the difference between decentralized and centralized methods is only significant in the large-batch regime. By integrating the findings of Theorem III.8 and Corollaries III.9 and IV.1, we deduce that network heterogeneity and graph structure inject additional noisy terms into decentralized methods and these facilitate their escape from sharp minima compared to the centralized method. However, these added noisy terms incorporate some deterioration into the long-term optimization performance. Furthermore, although consensus exhibits higher escaping efficiency, this comes at the expense of reduced optimization performance. This observation reveals an intrinsic trade-off between flatness and optimization within the context of multi-agent learning.

### v2

@@ III-A Modeling conditions
We introduce the following commonly-used assumptions.
First, to analyze how an algorithm escapes from a given location, it is essential to assume that the algorithm is operating near that point. This point of view is widely adopted in the literature when examining the ability of algorithms to escape saddle points [16, 20, 52] or local minima [33, 38, 47, 48]. For example, references [16, 20, 52] assume that the stochastic gradient algorithm begins from a point sufficiently close to a saddle point, where the gradient norm is small. The work [33] assumes the algorithm starts precisely at a local minimum, while [38] assumes the model is near a local minimizer, thus allowing the Hessian at the current point to be well approximated by the Hessian at the local minimizer. Such assumptions enable a focused analysis of the algorithms’ behavior near critical points. In our study, since we focus on analyzing the behavior of algorithms in the vicinity of a local minimizer, $w^{\star}$, we assume that all agents start from points close to it.
@@
(Starting points of all agents). All models initiate their updates from points sufficiently close to $w^{\star}$, namely,
@@
By applying Jensen’s inequality to (11), we obtain:
@@
Assumption III.1 is justified from traditional convergence results. For instance, assume we pick any step size $\mu^{\prime}\leq o(\mu)$, such as $\mu^{\prime}=\mu^{2}$. Then, it is well-known that the centralized stochastic gradient algorithm approaches an $O(\frac{\mu^{\prime}}{B})$ neighborhood of a local minimum of $J(w)$ after sufficient iterations in nonconvex environments (see, e.g., [7, 52, 20, 53]), namely,
$\lim_{n\to\infty}\mathds{E}\|\nabla J(\boldsymbol{w}_{k,n})-\nabla J(w^{\star})\|^{2}\leq O\left(\frac{\mu^{\prime}}{B}\right)=o\left(\frac{\mu}{B}\right)$ (13)
where $\nabla J(w^{\star})=0$ since $w^{\star}$ admits a local minimum of $J(w)$. Additionally, under the centralized setting, the influence of the term related to the difference between local models and their centroid is eliminated. Moreover, using the local convexity around $w^{\star}$, it follows from [6] that:
$\mathds{E}\|\boldsymbol{w}_{k,n}-w^{\star}\|^{2}\leq o\left(\frac{\mu}{B}\right),\;\mathds{E}\|\boldsymbol{w}_{k,n}-w^{\star}\|^{4}\leq o\left(\frac{\mu^{2}}{B^{2}}\right)$ (14)
where the term related to the network heterogeneity vanishes in the centralized setting. As a result, the bounds in Eqs. (13) and (14) are determined solely by gradient noise. In summary, with sufficiently small step sizes, agents can attain an $o(\frac{\mu}{B})$-neighborhood of $w^{\star}$.
@@
$\bar{H}\overset{\Delta}{=}\frac{1}{K}\sum_{k=1}^{K}\nabla^{2}J_{k}(w^{\star})=\frac{1}{K}\sum_{k=1}^{K}H_{k}^{\star}$ (17)
@@ III-B Properties of gradient noise
which corresponds to the case of $B=1$ (i.e., batch size equal to 1). In the mini-batch case, the gradient noise is instead given by:
@@
For later use, we define the gradient covariance matrix of the global risk function at $w^{\star}$ for the case $B=1$ as:
@@ III-C Network performance analysis
This approximation naturally raises the following question: How accurate can the short-term model in (47) approximate the true recursion in (41)? A similar issue has been addressed for nonconvex optimization problems in the vicinity of saddle points in [16]. Motivated by this analysis, we answer the above question in the following two lemmas, where we separately show the results for the decentralized and centralized methods over a finite time horizon.
Lemma III.5.
@@
See Appendices B, C and D. ∎
Lemma III.6.
(Deviation bounds of the centralized method). Under the same conditions of Lemma III.5, the second-order and fourth-order moments of ${\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}$, the second-order moment of ${\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}$, and the approximation error of the short term model, related to the centralized method are guaranteed to be upper bounded in a finite number of iterations $n\leq O(\frac{1}{\mu})$. Basically, it holds that
@@
$\Big{|}\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}\Big{|}\leq O(\mu^{1.5(1+\eta))})$ (56)
@@
See Appendices B, C and D. ∎
Comparing the bounds in Lemmas III.5 and III.6, we observe that the upper bounds on the second-order moments associated with the decentralized methods incorporate extra $O(\mu^{2})$ terms compared to the centralized method. These extra terms arise from the consensus distance, as defined in [27], which measures the distance between the local models and their centroid. This consensus distance is also related to both the network heterogeneity and the graph structure. Detailed proofs of these results can be found in Appendix B. We will later discuss how the presence of the $O(\mu^{2})$ terms impacts the escaping efficiency.
Lemmas III.5 and III.6 demonstrate that $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}$ and $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}\|^{2}$ dominate $|\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}|$ for all methods when $\mu$ is sufficiently small. In other words, the approximation error between the mean square deviation of the short-term and true models can be omitted compared with the size of the true models. Furthermore, we can manipulate the bounds in Lemmas III.5 and III.6 to obtain expression for $\mathrm{ER}_{n}$ performance. Using (7) and (9), we can approximate the $\mathrm{ER}_{n}$ as follows:
@@
which means that $\mathrm{ER}_{n}$ can be evaluated by means of the short-term recursion. Thus, the original recursion in (41) can be replaced by (47) in the small step-size regime. More details about the equality (III-C) can be found in Appendix E.
Building upon the results of Lemmas III.5 and III.6, we next analyze the escaping efficiency of the three algorithms by using (47), and establish the following theorem.
Theorem III.7.
@@
$e(n)=\frac{1}{4}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{R}_{s}\right)$ (62)
@@
See Appendix E. ∎
Theorem III.7 shows the escaping efficiency of the three algorithms around a local minimum over a finite time horizon. We emphasize that the $O(\frac{\mu}{B})$ or $O(\frac{\mu}{B})+O(\mu^{2})$ terms in (59)–(61) are the dominant contributors. Therefore, focusing on these terms is sufficient to compare the escaping efficiency of the three distributed algorithms. Basically, for all of three methods, larger $\mu$ enables higher escaping efficiency. Also, as mentioned in Appendix E, for the algorithms to successfully exit the local basin they would require $O(1/{\mu})$ iterations. Thus, if the algorithms cannot leave the local minimum in $O(1/\mu)$ iterations, then we say they cannot effectively escape the local basin and are therefore trapped in it. Furthermore, as discussed in Appendix E, decentralized methods implicitly include the additional $O(\mu^{2})$ terms related to the consensus distance. The network heterogeneity and graph structure ensure that the consensus distance is non-zero so that improve the escaping efficiency. That is, for homogeneous networks where all agents share same local minimizers $w^{\star}$, we have
@@
However, the difference between the centralized and decentralized methods is only significant in the large-batch regime. Specifically, if $B$ is not sufficiently large such that $\eta<1$, then we have $21+\eta<$ 2 under which the additional $O(\mu^{2})$ terms in (60) and (61) will be dominated by the $o(\mu^{1+\eta})$ terms associated with the approximated error due to (47) and the initialization condition in Assumption III.1. As the batch size $B$ increases, however, the effect of the gradient noise will progressively decrease and the influence of the $O(\mu^{2})$ terms will become more prominent. In the extreme scenario when the full-batch gradient descent is applied, the centralized method becomes completely noise-free, which means that there is no noise to help the algorithm escape from local minima. In contrast, the noisy terms related to network heterogeneity and graph structure continue to exist in decentralized approaches. This makes decentralized methods more effective at escaping from local minima than the centralized method in the large-batch regime. However, larger values of $\mathrm{ER}_{n}$, while good for escaping efficiency, they nevertheless worsen the optimization performance. We will elaborate on this trade-off later in the steady state when $n\to\infty$.
@@
Corollary III.8.
(Smaller $\mathrm{ER}_{n}$ for diffusion). Under the same conditions of Theorem III.7, it holds that
@@
See Appendix F. ∎
Corollary III.8 implies that consensus runs farther away from the local minimum than the diffusion strategy for the same number of iterations due to its worse excess-risk performance. This translates into faster escape for consensus compared to diffusion.
@@ IV Trade-off between flatness and optimization
Suppose we run a decentralized or centralized algorithm to solve a possibly nonconvex optimization problem. One useful question is to investigate where the algorithm would prefer to go if it escapes from the current basin. In other words, assume the algorithm escapes from the basin around some local minima and starts evolving until it settles down in the basin of another local minimum, we would like to examine what preferential properties does this second minimum have relative to the earlier one. To answer this question, we will relate the escaping efficiency of algorithms measured by $\mathrm{ER}_{n}$ and the flatness of local minima measured by $\mathrm{Tr}(\bar{H})$.
@@
$e(n)=$ $\frac{1}{4}\mathrm{Tr}\bigg{(}\Big{(}I-(I-\mu\bar{H})^{2(n+1)}\Big{)}{\bar{R}_{s}}\bigg{)}$
$\leq$ ${\frac{1}{4}\lambda_{\max}(\bar{R}_{s})\mathrm{Tr}\Big{(}I-(I-\mu\bar{H})^{2(n+1)}\Big{)}}$
$\leq$ ${\frac{1}{4}\lambda_{\max}(\bar{R}_{s})\sum\limits_{i=1}^{M}\left(1-(1-\mu\lambda_{i}(\bar{H}))^{2(n+1)}\right)}$
=$\overset{\Delta}{=}$ $\mathcal{U}\big{(}e(n)\big{)}$ (68)
where $\lambda_{i}(\bar{H})$ represents the $i_{th}$ eigenvalue of $\bar{H}$, and $\lambda_{\max}(\bar{R}_{s})$ denotes the largest eigenvalue of $\bar{R}_{s}$. Also,
@@
Observe that these upper bounds on the escaping efficiency are positively correlated with the eigenvalues of $\bar{H}$, i.e., the sharpness of the local minimum. Substituting (71)–(73) into (67) separately, we find that flat minima where $\mathrm{Tr}(\bar{H})$ is small provide high probability guarantees that the algorithms stay around the corresponding local basin.
Combining the discussion of both one and higher-dimensional cases, we conclude that the three algorithms prefer to stay around flat minima, which means that they favor flatter basin if they successfully escape from a current minimum. Moreover, as discussed earlier for Theorem III.7, higher escaping efficiency makes decentralized algorithms more likely to escape from a local minimum than the centralized method. This fact further indicates that decentralized methods favor flatter minima than their centralized counterpart. Incidentally, the expression for $\mathrm{ER}_{n,cen}$ in Theorem III.7 is consistent with the well-known observations in the literature [29, 32], which highlight that larger step sizes or smaller batch sizes with the mini-batch SGA in the single-agent setup steer the model toward flatter minima. However, as already noted before, higher escaping efficiency may deteriorate the optimization performance. This motivates us to analyze the excess-risk performance in the long run when $n\to\infty$, which corresponds to the optimization performance in the steady state.
@@
(Steady-state excess risks). Consider a network of agents running distributed algorithms covered by (47). After sufficient iterations such that $n\to\infty$, under assumptions II.1, III.1, III.2 and III.3, and assuming the algorithms are already trapped in the basin of a local minimizer $w^{\star}$ and $J(w)$ is locally strongly-convex around $w^{\star}$, it holds that
$\mathrm{ER}_{\infty,cen}=$ $\frac{\mu}{4B}\mathrm{Tr}\left(\bar{R}_{s}\right)\pm O\left(\mu^{1.5(1+\eta)}\right)$ (74)
$\mathrm{ER}_{\infty,con}=$ $\frac{\mu}{4B}\mathrm{Tr}\left(\bar{R}_{s}\right)+\frac{\mu^{2}}{2K}\|d^{\sf T}\mathcal{V}_{\alpha}(I-\mathcal{P}_{\alpha})^{-1}\|^{2}_{I\otimes\bar{H}}$
@@
$\mathrm{ER}_{\infty,dif}=$ $\frac{\mu}{4B}\mathrm{Tr}\left(\bar{R}_{s}\right)+\frac{\mu^{2}}{2K}\|d^{\sf T}\mathcal{V}_{\alpha}P_{\alpha}(I-\mathcal{P}_{\alpha})^{-1}\|^{2}_{I\otimes\bar{H}}$
@@
See Appendix G. ∎
In Corollary IV.1, the factors that help algorithms escape local minima are now seen to adversely affect the optimization performance. Again, the difference between decentralized and centralized methods is only significant in the large-batch regime. By integrating the findings of Theorem III.7 and Corollaries III.8 and IV.1, we deduce that network heterogeneity and graph structure inject additional noisy terms into decentralized methods and these facilitate their escape from sharp minima compared to the centralized method. However, these added noisy terms incorporate some deterioration into the long-term optimization performance. Furthermore, although consensus exhibits higher escaping efficiency than diffusion, this comes at the expense of reduced optimization performance. This observation reveals an intrinsic trade-off between flatness and optimization within the context of multi-agent learning. In addition, the results indicate a trade-off in the choice of the step size $\mu$. While larger $\mu$ can help the algorithms escape from sharp minima, it also increases the risk that the algorithms may fail to converge altogether. This motivates the use of the piecewise learning-rate schedule [55, 56] in practice. That is, a moderately small $\mu$ is applied during the initial stage of training to assist in escaping sharp minima. Later, a smaller $\mu$ is adopted to ensure convergence.